OpenStudy (anonymous):

(Stoichiometry) The annual production of sulfur dioxide from burning coal and fossil fuels, for auto exhaust, and from other activities is about 26 million tons. How much sulfur, present in the original materials, would result in that quantity of SO2? S+O2 --> SO2

OpenStudy (anonymous):

Let x be the amount of sulfur in million tonnes. So the mass of O2 = 26-x The composition must be same throughout the compound. So, (26-x/x) = (32/32) , 26-x = x , x= 13 million tonnes. Therefore 13 million tonnes of sulfur is present.

OpenStudy (anonymous):

still can't understand. can you explain it further? thank you.

OpenStudy (australopithecus):

We have 26 million tons of SO2 we want to figure out how much Sulfur we need to acquire 26million tons of SO2 assuming O2 is in excess we can just use gravimeter factor to solve this, \[Mass_{SO_2}*\frac{Molecular Mass_S}{Molecular Mass_{SO_2}} = Mass_S\] so we have \[26*\frac{(32.07g/mol)}{(32.07g/mol+15.9994g/mol+15.9994g/mol)} = 13\] thus we have 13million tons of Sulpher Gravimeter factor is pretty easy to use, if you have any questions about it feel free to ask

OpenStudy (australopithecus):

One of the key things you should notice is that the g/mol units cancel out and you are just left with mass units, this is essentially giving us the ratio of the mass that one element takes up in that given compound

OpenStudy (australopithecus):

sorry that is gravimetric factor NOT gravimeter factor, I blame spell check